### v1

@@ Abstract
The IC3 algorithm, also known as PDR, is a SAT-based model checking algorithm that has significantly influenced the field in recent years due to its efficiency, scalability, and completeness. It utilizes SAT solvers to solve a series of SAT queries associated with relative induction. Based on our observations of the unique characteristics of SAT queries in IC3, this paper introduces GipSAT, a lightweight SAT solver specifically optimized for IC3. By observing that SAT queries do not necessarily require decisions on all variables, GipSAT calculates a subset of variables that need to be decided before each solving, while ensuring that the result remains unaffected. By observing that the overhead of binary heap operations in VSIDS is not negligible, GipSAT utilizes buckets instead of binary heap to achieve constant-time operations. GipSAT supports temporary clauses without the need to allocate a new activation variable before each solving, thus eliminating the need to reset solvers. The comprehensive evaluation demonstrates a significant performance improvement achieved by GipSAT. When compared to Minisat, GipSAT achieves an average speedup of $3.61$ times in solving time.
@@ 1 Introduction
Model checking [15, 16] is a powerful formal verification technique widely used in system design. Given a transition system and a property describing the desired system behavior, it can efficiently and automatically detect bugs or prove system safety.
@@
IC3 works by constructing SAT problems regarding the reachable states of the model being verified. With the help of a SAT solver, solving these problems yields answers that facilitate the verification process. Therefore, currently, almost all implementations of IC3 essentially have a two-layer structure: the top layer is the IC3 algorithm layer, which drives the verification process by posing a series of SAT queries, and the bottom layer is the SAT solver layer, which handles the queries posed by the algorithm layer.
@@
In this paper, we present a lightweight SAT solver, GipSAT, which is deeply optimized for the IC3 algorithm based on several observations, enabling it to handle SAT queries in IC3 more efficiently. The contributions of this paper are summarized as follows:
@@
  • We introduce a novel data structure that allows VSIDS operations to be performed in constant time.
  • We propose a method that eliminates the need to allocate an activation variable before each solving, thereby avoiding SAT solver resetting.
  • We conduct a comprehensive performance evaluation of GipSAT. The experimental results show that GipSAT significantly improves the efficiency and scalability of the IC3 algorithm.
Organization: In Section 2, we provide some background. The motivations behind our work are detailed in Section 3. The design and implementation of GipSAT are described in Section 4. A comprehensive evaluation of GipSAT on various IC3 implementations is presented in Section 5. Section 6 discusses related works. Finally, in Section 7, we draw conclusions and outline future directions.
@@ 2.1 Basics and Notations
We use notations such as $x,y$ for Boolean variables, and $X,Y$ for sets of Boolean variables. The terms $x$ and $\lnot x$ are referred to as literals. A cube is a conjunction of literals, while a clause is a disjunction of literals. A Boolean formula in Conjunctive Normal Form (CNF) is a conjunction of clauses. It is often convenient to treat a clause or a cube as a set of literals, and a CNF as a set of clauses. For instance, given a CNF formula $F$, a clause $c$, and a literal $l$, we write $l\in c$ to indicate that $l$ occurs in $c$, and $c\in F$ to indicate that $c$ occurs in $F$. A formula $F$ implies another formula $G$, if every satisfying assignment of $F$ satisfies $G$, denoted as $F\Rightarrow G$ or $F\models G$. We use $\mathcal{V}(F)$ to represent the set of all variables appeared in formula $F$.
A transition system, denoted as $S$, can be defined as a tuple $\langle X,Y,I,T\rangle$. Here, $X$ and $X^{\prime}$ represent the sets of state variables for the current state and the next state, respectively, while $Y$ represents the set of input variables. The Boolean formula $I(X)$ represents the initial states, and $T(X,Y,X^{\prime})$ describes the transition relation. State $s_{2}$ is a successor of state $s_{1}$ if and only if $(s_{1},s_{2}^{\prime})$ is an assignment of $T$ ($(s_{1},s_{2}^{\prime})\models T$). A safety property $P(X)$ is a Boolean formula over $X$. A system $S$ satisfies $P$ if and only if all reachable states of $S$ satisfy $P$.
@@ 2.3 IC3 Algorithm
Algorithm 1 provides an overview of the IC3 algorithm. This algorithm incrementally constructs frames by iteratively performing the blocking phase and the propagation phase. During the blocking phase, the IC3 algorithm focuses on making $F_{k}\Rightarrow P$. It iteratively get a cube $c$ such that $c\models\lnot P$, and block it recursively. This process involves attempting to block the cube’s predecessors if it cannot be blocked directly. It continues until the initial states cannot be blocked, indicating that $\lnot P$ can be reached from the initial states in $k$ transitions, which violates the property. In cases where a cube can be confirmed as blocked, IC3 proceeds to expand the set of blocked states through a process called generalization. This involves dropping literals and ensuring that the resulting clause remains relatively inductive, with the objective of obtaining a minimal inductive clause. The propagation phase tries to push lemmas to the top frame. If a lemma $c$ in $F_{i}\setminus F_{i+1}$ is also inductive relative to $F_{i}$, then push it into $F_{i+1}$. During this process, if two consecutive frames become identical ($F_{i}=F_{i+1}$), then the inductive invariant is found and the safety of this model can be proved.
@@ 3 Motivations
We analyzed the characteristics of SAT queries in Minisat based rIC3 [9]. The analysis was conducted on the complete benchmark of the 2020 and 2024 Hardware Model Checking Competition (HWMCC). For a detailed description of the experimental setup, refer to Section 5. The results are presented in Table 1. It can be observed that the SAT queries generated by IC3 exhibit unique characteristics: they are solved quickly, with an average solving time of less than 1ms, but the number of queries is very large. These SAT queries differ significantly from those encountered in SAT competitions.
We aim to implement a lightweight SAT solver to handle a large number of relatively simple problems more efficiently, thereby improving overall performance. Additionally, we have identified certain characteristics in solving these SAT problems and plan to further optimize the SAT solver for IC3 based on these observations.
  • Observation 1: $Q_{relind}$: $sat(F_{i}\land c\land T\land\lnot c^{\prime})$ represents whether there exists a predecessor of any state in $\lnot c$ (where $\lnot c$ is a cube, representing a set of states) that satisfies $F_{i}$ but is not in $\lnot c$. We have observed that solving $Q_{relind}$ does not require considering assignments for all variables. Fig. 1 shows an AIG model. The values of variables in $c^{\prime}$ are determined solely by $COI(c^{\prime})$, as illustrated by the shaded region. The values of variables outside the shaded region do not affect the result. From the last column of Table 1, it can be observed that the $c$ in $Q_{gen}$ and $Q_{push}$ is typically very small in size. As a result, the size of $COI(c^{\prime})$ is considerably small compared to the total number of variables. We also calculated the average percentage of variables in $COI(c^{\prime})$ for all $Q_{relind}$ relative to the total number of variables, as shown in Fig. 2. It can be observed that, in most cases, the number of variables that need to be determined is only around 20% of the total number of variables.
@@
    Figure 2: The average percentage of the number of variables of $COI(c^{\prime})$ relative to the total number of variables of $Q_{relind}$ for each case.
    Figure 3: The overhead of VSIDS operations for each case.
@@ 4.1 Decide and Propagate in Domain
Observation 1 shows that it is unnecessary to consider all variables when solving $Q_{relind}$. For the satisfiability of the formula $F_{i}\land c$, we only need to consider the variables that appear in it, which are $\mathcal{V}(F_{i})\cup\mathcal{V}(c)$. Similarly, for the satisfiability of $T\land\lnot c^{\prime}$, the values of the variables in $c^{\prime}$ are determined solely by the $COI(c^{\prime})$, as illustrated by the shaded region in Fig. 1. Therefore, only the variables in $COI(c^{\prime})$ need to be considered. This significantly reduces the number of variables that the SAT solver needs to decide and assign during BCP.
We refer to the set of variables that need to be considered during solving as the domain. For solving $Q_{relind}$, it is sufficient to consider $\mathcal{V}(F_{i})\cup\mathcal{V}(c)\cup COI(c^{\prime})$ as the domain, rather than all variables. GipSAT maintains a data structure for $\mathcal{V}(F_{i})$, which is updated whenever a new lemma is added. Before each solving step, GipSAT computes $COI(c^{\prime})$ by recursively traversing the fanins. At the start of the solving process, GipSAT pushes all variables of the current domain into VSIDS’s decision-required set, ensuring that all variables in the domain will be decided. During Boolean constraint propagation, if a clause includes a variable not in the current domain, GipSAT skips this clause and proceeds to the next one.
It can also be noticed that the difference between $c^{\prime}$ in every two adjacent $Q_{gen}$ is only one literal. This suggests that calculating $COI(c^{\prime})$ for every $Q_{gen}$ is unnecessary. Instead, GipSAT performs this calculation initially during generalization and upon successful drops, as depicted in Algorithm 2. GipSAT recalculates it upon successful drops because sometimes the unsat core can drop multiple variables at once compared to before.
@@ 4.2 VSIDS by Bucket
Observation 2 illustrates that the overhead of VSIDS is non-negligible. Mitigating this overhead could potentially lead to improved overall performance. Inspired by bucket sorting, GipSAT introduces a novel data structure within VSIDS, enabling both push and pop operations to be performed in constant time. GipSAT predefines a fixed number of buckets (defaulting to 15) and assigns each variable to a specific bucket. This design ensures that variables in lower-numbered buckets have higher VSIDS scores than those in higher-numbered buckets. To maintain this guarantee amid dynamic changes in variable scores, GipSAT utilizes an interval heap for each bucket. The interval heap is a double-ended priority queue that can efficiently pop either the maximum or minimum value in logarithmic time [26]. When the score of variable $v$ increases, the $update$ function in Algorithm 3 is invoked. This function first locates the bucket $b$ where $v$ resides. Then, it checks whether the maximum score in bucket $b$ is greater than the minimum score in bucket $b-1$. If so, it pops the variable with the maximum score from bucket $b$ and pushes it into bucket $b-1$. Next, it pops the variable with the minimum score from bucket $b-1$ and pushes it into bucket $b$. This process is repeated on bucket $b-1$ to ensure the monotonicity of scores across buckets. GipSAT creates a queue for each bucket, and during the variable decision, it selects a variable from the non-empty queue of the smallest bucket, as described in the $pop$ function of Algorithm 3. During backtracking, it pushes the unassigned variables into their respective queues based on their buckets, as described in the $push$ function. This method effectively reduces the operations of VSIDS from logarithmic time to constant time. However, it also compromises the accuracy of VSIDS. Nevertheless, since the queries are relatively simple, it may not necessitate very high accuracy.
@@ 4.3 Without Resetting
Observation 3 suggests that the SAT solver needs to be reset periodically because of the activation variables. However, it also clears all learned clauses and heuristic scores. GipSAT does not require resetting. It can be noticed that all the learnt clauses derived from the temporary clause contain the activation variable [13]. GipSAT allocates a fixed variable in advance and uses this as the activation variable in every query. If a learnt clause includes this variable, GipSAT records it in a special vector. After each solving, GipSAT deletes all learnt clauses stored in this special vector from the clause database, along with the temporary clause itself. Consequently, the temporary clause and all its derived learned clauses are removed, ensuring that the activation variable is not included in any clause. This makes the variable free again, available for reuse in the next solving. This approach requires only one additional variable, thus eliminating the need to reset the SAT solver.
4.4 Implementation
Instead of implementing these optimizations on an existing SAT solver, we developed a new lightweight CDCL-based SAT solver from scratch, called GipSAT. It is implemented in Rust, a modern programming language designed to offer both performance and safety. Additionally, It provides a C++ interface, making it compatible with C++ projects.

### v2

@@ Abstract
The IC3 algorithm, also known as PDR, is a SAT-based model checking algorithm that has significantly influenced the field in recent years due to its efficiency, scalability, and completeness. It utilizes SAT solvers to solve a series of SAT queries associated with relative induction. In this paper, we introduce several optimizations for the SAT solver in IC3 based on our observations of the unique characteristics of these SAT queries. By observing that SAT queries do not necessarily require decisions on all variables, we compute a subset of variables that need to be decided before each solving process while ensuring that the result remains unaffected. Additionally, noting that the overhead of binary heap operations in VSIDS is non-negligible, we replace the binary heap with buckets to achieve constant-time operations. Furthermore, we support temporary clauses without the need to allocate a new activation variable for each solving process, thereby eliminating the need to reset solvers. We developed a novel lightweight CDCL SAT solver, GipSAT, which integrates these optimizations. A comprehensive evaluation highlights the performance improvements achieved by GipSAT. Specifically, the GipSAT-based IC3 demonstrates an average speedup of $3.61$ times in solving time compared to the IC3 implementation based on MiniSat.
@@ 1 Introduction
Model checking [16, 17] is a powerful formal verification technique widely used in system design. Given a transition system and a property describing the desired system behavior, it can efficiently and automatically detect bugs or prove system correctness.
@@
IC3 works by constructing SAT problems regarding the overapproximate reachable states of the model being verified. With the help of a SAT solver, solving these problems yields answers that facilitate the verification process. Therefore, currently, almost all implementations of IC3 essentially have a two-layer structure: the top layer is the IC3 algorithm layer, which drives the verification process by posing a series of SAT queries, and the bottom layer is the SAT solver layer, which handles the queries posed by the algorithm layer.
@@
In this paper, we propose several optimizations for the SAT solver used in IC3, based on our analysis of the unique characteristics of SAT queries. We introduce a lightweight SAT solver, GipSAT, which incorporates these optimizations to enhance its efficiency in processing SAT queries within IC3. The key contributions of this paper are summarized as follows:
@@
  • We introduce a novel data structure that replaces the binary heap, enabling VSIDS operations to be performed in constant time.
  • We propose a method that supports temporary clauses by reusing the same activation variable instead of allocating a new one before each solving process, thus eliminating the need for SAT solver resetting.
  • We introduce a novel lightweight SAT solver, GipSAT, which incorporates these optimizations. We provide a detailed interface for GipSAT.
  • We conduct a comprehensive performance evaluation of GipSAT. The experimental results demonstrate that GipSAT significantly enhances the efficiency and scalability of the IC3 algorithm.
@@ 2.1 Basics and Notations
We use notations such as $x,y$ for Boolean variables, and $X,Y$ for sets of Boolean variables. The terms $x$ and $\lnot x$ are referred to as literals. A cube is a conjunction of literals, while a clause is a disjunction of literals. A Boolean formula in Conjunctive Normal Form (CNF) is a conjunction of clauses. It is often convenient to treat a clause or a cube as a set of literals, and a CNF as a set of clauses. For instance, given a CNF formula $F$, a clause $c$, and a literal $l$, we write $l\in c$ to indicate that $l$ occurs in $c$, and $c\in F$ to indicate that $c$ occurs in $F$. A formula $F$ implies another formula $G$, if every satisfying assignment of $F$ satisfies $G$, denoted as $F\Rightarrow G$. We use $\mathcal{V}(F)$ to represent the set of all variables appeared in formula $F$.
A transition system, denoted as $S$, can be defined as a tuple $\langle X,Y,I,T\rangle$. Here, $X$ and $X^{\prime}$ represent the sets of state variables for the current state and the next state, respectively, while $Y$ represents the set of input variables. The Boolean formula $I(X)$ represents the initial states, and $T(X,Y,X^{\prime})$ describes the transition relation. State $s_{2}$ is a successor of state $s_{1}$ if and only if $(s_{1},s_{2}^{\prime})\Rightarrow T$. A safety property $P(X)$ is a Boolean formula over $X$. A system $S$ satisfies $P$ if and only if all reachable states of $S$ satisfy $P$.
@@ 2.3 IC3 Algorithm
Algorithm 1 provides an overview of the IC3 algorithm. This algorithm incrementally constructs frames by iteratively performing the blocking phase and the propagation phase. During the blocking phase, the IC3 algorithm focuses on making $F_{k}\Rightarrow P$. It iteratively get a cube $c$ such that $c\Rightarrow\lnot P$, and block it recursively. This process involves attempting to block the cube’s predecessors if it cannot be blocked directly. It continues until the initial states cannot be blocked, indicating that $\lnot P$ can be reached from the initial states in $k$ transitions, which violates the property. In cases where a cube can be confirmed as blocked, IC3 proceeds to expand the set of blocked states through a process called generalization. This involves dropping literals and ensuring that the resulting clause remains relatively inductive, with the objective of obtaining a minimal inductive clause. The propagation phase tries to push lemmas to the top frame. If a lemma $c$ in $F_{i}\setminus F_{i+1}$ is also inductive relative to $F_{i}$, then push it into $F_{i+1}$. During this process, if two consecutive frames become identical ($F_{i}=F_{i+1}$), then the inductive invariant is found and the safety of this model can be proved.
@@ 3 Motivations
We analyzed the characteristics of SAT queries in Minisat-based rIC3 [9]. The analysis was conducted on the complete benchmarks of the 2020 and 2024 Hardware Model Checking Competition (HWMCC). For a detailed description of the experimental setup, please refer to Section 6. The results are presented in Table 1. It can be observed that the SAT queries generated by IC3 exhibit unique characteristics: they are solved quickly, with an average solving time of less than 1ms, but the number of queries is very large. Additionally, we have the following observations:
  • Observation 1: $Q_{relind}$: $sat(F_{i}\land c\land T\land\lnot c^{\prime})$ represents whether there exists a predecessor of any state in $\lnot c$ (where $\lnot c$ is a cube, representing a set of states) that satisfies $F_{i}$ but is not in $\lnot c$. We have observed that solving $Q_{relind}$ does not require considering assignments for all variables. Fig. 1 shows an AIG model. The values of variables in $c^{\prime}$ are determined solely by $COI(c^{\prime})$, as illustrated by the shaded region. The values of variables outside the shaded region do not affect the result. From the last column of Table 1, it can be observed that the $c$ in $Q_{gen}$ and $Q_{push}$ is typically very small in size. Consequently, the size of $COI(c^{\prime})$ is likely to be small compared to the total number of variables. We also calculated the average percentage of variables in $COI(c^{\prime})$ for all $Q_{relind}$ relative to the total number of variables, as shown in Fig. 2. It can be observed that, in most cases, the number of variables that need to be determined is only around 20% of the total number of variables. We may be able to use this information to reduce the number of variables that need to be assigned in each solving iteration. However, regular SAT solvers can only obtain the CNF representation of the AIG and cannot capture the dependency relationships between variables.
@@
    Figure 2: The number of cases (x-axis) in which the average percentage of variables in $COI(c^{\prime})$ relative to the total number of variables in $Q_{relind}$ exceeds a given value (y-axis).
    Figure 3: The number of cases (x-axis) in which the overhead of VSIDS operations exceeds a given value (y-axis).
@@ 4.1 Decide and Propagate in Domain
Observation 1 shows that it is unnecessary to consider all variables when solving $Q_{relind}$. We analyze which variables need to be considered through the following theorems. Considering the formula $Q_{relind}$: $sat(F_{i}\land c\land T\land\lnot c^{\prime})$, we can analyze the subset of variables that need to be taken into account without affecting the satisfiability. To determine the satisfiability of $F_{i}\land c$, we only need to consider the variables that appear in this formula, which are $\mathcal{V}(F_{i})\cup\mathcal{V}(c)$. Similarly, for the satisfiability of $T\land\lnot c^{\prime}$, the values of the variables in $\lnot c^{\prime}$ are solely determined by the $COI(c^{\prime})$. Therefore only the variables in $COI(c^{\prime})$ need to be considered. As a result, by considering only the variables in $\mathcal{V}(F_{i})\cup\mathcal{V}(c)\cup COI(c^{\prime})$, we can determine the satisfiability of the formula $Q_{relind}$.
We refer to the set of variables that need to be considered during solving as the domain. For solving $Q_{relind}$, it is sufficient to use $\mathcal{V}(F_{i})\cup\mathcal{V}(c)\cup COI(c^{\prime})$ as the domain, instead of all variables. This significantly reduces the number of variables that the SAT solver needs to decide and assign during BCP. We need to efficiently compute the domain before each solving process. For $\mathcal{V}(F_{i})$, we maintain a data structure within the SAT solver that tracks and updates it whenever a new lemma is added. For $\mathcal{V}(c)$, we simply extract all the variables that appear in $c$. Finally, $COI(c^{\prime})$ is computed by recursively traversing the fanins of $c^{\prime}$.
At the beginning of the solving process, we push all variables from the computed domain into the VSIDS decision-required set, ensuring that all domain variables will be decided. During BCP, if a clause includes a variable not in the current domain, we skip that clause and move on to the next one. This approach ensures that variables outside the domain are not assigned during BCP.
It can also be noticed that the difference between $c^{\prime}$ in every two adjacent $Q_{gen}$ is only one literal. This suggests that calculating $COI(c^{\prime})$ for every $Q_{gen}$ is unnecessary. Instead, We perform this calculation initially during generalization and upon successful drops, as depicted in Algorithm 2. We recalculate it upon successful drops because sometimes the unsat core can drop multiple variables at once compared to before.
@@ 4.2 VSIDS by Bucket
Observation 2 highlights that the overhead of VSIDS is non-negligible, and reducing this overhead could potentially enhance overall performance. Inspired by bucket sorting, we introduce a novel data structure for VSIDS, that enables both push and pop operations to be performed in constant time. We predefine a fixed number of buckets (defaulting to 15) and assign each variable to a specific bucket. This design ensures that variables in lower-numbered buckets have higher VSIDS scores than those in higher-numbered buckets.
To maintain this guarantee amid dynamic changes in variable scores, we utilize an interval heap for each bucket. The interval heap is a double-ended priority queue that can efficiently pop either the maximum or minimum value in logarithmic time [27]. Each bucket’s interval heap stores all the variables belonging to that bucket, using their scores as the sorting criterion. When the score of variable $v$ increases, the $update$ function in Algorithm 3 is invoked. This function first locates the bucket $b$ where $v$ resides. Then, it checks whether the maximum score in bucket $b$ is greater than the minimum score in bucket $b-1$. If so, it pops the variable with the maximum score from bucket $b$ and pushes it into bucket $b-1$. Next, it pops the variable with the minimum score from bucket $b-1$ and pushes it into bucket $b$. This process is repeated on bucket $b-1$ to ensure the monotonicity of scores across buckets.
The interval heap is used to maintain the bucket corresponding to each variable, with changes occurring only when the score is updated. Additionally, we create a queue for each bucket to store the variables waiting to be decided within that bucket. The variables in the queue are unordered. During variable decision, the $pop$ function in Algorithm 3 selects a variable from the non-empty queue of the smallest bucket. During backtracking, the $push$ function places unassigned variables back into their respective queues based on their buckets.
This approach retains logarithmic time complexity for score updates while reducing the time complexity of the push and pop operations in VSIDS from logarithmic to constant time. However, this improvement comes at the cost of reduced accuracy in VSIDS, as the variable selected each time is from the bucket with the highest score but not necessarily the highest-scoring variable within that bucket. Nonetheless, since the queries are relatively simple, very high accuracy may not be critical.
@@ 4.3 Without Resetting
Observation 3 indicates that the SAT solver requires periodic resets due to the activation variables. However, this process also results in the loss of all learned clauses and heuristic scores. The key to supporting temporary clauses is to identify and remove all learned clauses that are directly or indirectly derived from temporary clauses after each solving process. We establish the following theorem:
Theorem 4.1
To solve $Q_{relind}$ with a temporary clause $\mathbf{c}$ by utilizing the activation variable $\mathbf{a}$, we consider the formula: $\text{sat}(F_{i}\land(c\lor a)\land T\land\lnot c^{\prime}\land\lnot a)$, where $\lnot c^{\prime}\land\lnot a$ serves as the temporary assumption. If a learned clause $\mathbf{lc}$ is derived from $c\lor a$, then $\mathbf{lc}$ must contain the literal $\mathbf{a}$.
Proof
In a CDCL solver, every learned clause is the result of a sequence of resolution steps [13, 37]. We prove this theorem by mathematical induction. Initially, the only clause containing $a$ is $c\lor a$, and we assume that no clause contains the literal $\lnot a$. Any learned clause derived from $c\lor a$ is either a direct or indirect resolvent of $c\lor a$. Since we assume that no clause contains $\lnot a$, $a$ cannot be used as a pivot in resolution. Consequently, every resolution step involving $c\lor a$ must preserve the literal $a$, implying that all learned clauses derived from $c\lor a$ must contain $a$. Furthermore, clauses not derived from $c\lor a$ cannot contain either $a$ or $\lnot a$, ensuring that our assumption remains valid throughout the process.
Theorem 4.1 states that learned clauses derived from the temporary clause contain the activation variable. Therefore, if all learned clauses containing the activation variable are identified and removed after each solving process, the temporary clause can also be safely deleted.
We use only one activation variable to support the temporary clause. A fixed variable is allocated in advance and serves as the activation variable for every query. During conflict clause analysis, if the learned clause contains this variable, it is recorded in a special vector. After each solving process, all learned clauses stored in this vector are removed from the clause database, along with the temporary clause itself. As a result, both the temporary clause and all its derived learned clauses are eliminated, ensuring that the activation variable does not appear in any remaining clauses. This frees the variable, making it available for reuse in the next solving process. Since this approach requires only a single additional variable, it eliminates the need to reset the SAT solver.
5 Implementation
Instead of implementing these optimizations on an existing SAT solver, we developed a new lightweight CDCL-based SAT solver from scratch, called GipSAT. It is implemented in Rust, a modern programming language designed to offer both performance and safety.